# Phantom Crossing and Oscillating Dark Energy with $F(R)$ Gravity

### v1

@@ IX Conclusions
In this paper, we studied the possibility of having dynamical DE which crosses the phantom divide line and oscillates at late times. We examined this problem both theoretically and numerically and proved that $F(R)$ gravity provides a consistent phenomenological framework for naturally describing the DE era without resorting to phantom fields. Specifically, we discussed the inverse-phantom crossing in $F(R)$ gravity from a theoretical viewpoint and discussed the conditions that the $F(R)$ gravity fluid must satisfy in order to achieve the inverse phantom crossing. We also discussed the apparent phantom crossing issue, further extending the considerations made. Furthermore, we developed the theoretical study of several models that may exhibit phantom crossing and discussed from a theoretical viewpoint these models and their DE behaviour, in a qualitative way, though. Moreover, several oscillating forms of the Hubble rate were considered, in terms of the $e$-foldings number and their realisation in $F(R)$ gravity was also considered. In addition, we considered the effects of a scalaron as dark matter with a curvature-dependent mass on the DE era. We also studied in a quantitative way the predictions on the DE era of two explicit $F(R)$ gravity models, one of which is also compatible with the DESI data, apart from the Planck data. We studied the behaviour of the deceleration parameter and the DE EoS, and showed that it is possible to obtain a phantom crossing EoS and, in addition, an oscillating DE EoS at late times. Thus, we proved that indeed $F(R)$ gravity offers a consistent framework which can generate a phantom to quintessence transition in a natural way without phantom fields, and moreover, some of the models realise DE oscillations, again built in the model, not artificially added. Now the question is why should $F(R)$ gravity should be considered a viable extension of Einstein’s gravity. The answer is simple: Einstein’s gravitational Lagrangian density is described simply by $R$, so an Occam’s razor approach is that an $F(R)$ gravity extends it. Of course, someone may also seek if such oscillating and phantom divide crossing phenomena occur in other gravities, like Einstein-Gauss-Bonnet theories, but the natural extension of Einstein’s gravity is $F(R)$ gravity and it enjoys an elevated role among all modified gravity models.
@@ Acknowledgements
We are indebted to Taishi Katsuragawa for discussions. This work was partially supported by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M, Spain (S.D.O).

### v2

@@ IX Conclusions
In this paper we studied the possibility of having dynamical DE which crosses the phantom divide line and oscillates at late times. We examined this problem both theoretically and numerically and proved that $F(R)$ gravity provides a consistent phenomenological framework for naturally describing the DE era without resorting to phantom fields. Specifically, we discussed the inverse-phantom crossing in $F(R)$ gravity from a theoretical viewpoint and discussed the conditions that the $F(R)$ gravity fluid must satisfy in order to achieve the inverse phantom crossing. We also discussed the apparent phantom crossing issue, further extending the considerations made. Furthermore, we developed the theoretical study of several models that may exhibit phantom crossing and discussed from a theoretical viewpoint these models and their DE behavior, in a qualitative way though. Moreover, several oscillating forms of the Hubble rate were considered, in terms of the $e$-foldings number and their realization in $F(R)$ gravity was also considered. In addition, we studied the effects of a scalaron as dark matter with a curvature-dependent mass on the DE era. We also studied in a quantitative way the predictions on the DE era of two explicit $F(R)$ gravity models, one of which is also compatible with the DESI data apart from the Planck data. We considered the behavior of the deceleration parameter and the DE EoS and showed that it is possible to obtain a phantom crossing EoS and in addition an oscillating DE EoS at late times. Thus we proved that indeed $F(R)$ gravity offers a consistent framework which can generate a phantom to quintessence transition in a natural way without phantom fields and moreover some of the models realize DE oscillations, again built in the model, not artificially added. Now the question is why should $F(R)$ gravity should be considered a viable extension of Einstein’s gravity. The answer is simple, Einstein’s gravitational Lagrangian density is described simply by $R$, so an Occam’s razor approach is that an $F(R)$ gravity extends it. Of course, someone may also seek if such oscillating and phantom divide crossing phenomena occur in other gravities, like Einstein-Gauss-Bonnet theories, but the natural extension of Einstein’s gravity is $F(R)$ gravity and it enjoys an elevated role among all modified gravity models.
@@ Acknowledgements
This work was partially supported by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M, Spain (S.D.O).